OpenStudy (ttop0816):

Choose the polar form of the complex number -1 + i. sqrt2(cos45° + isin45°) sqrt2(cos135° + isin135°) 2(cos45° + isin45°) 2(cos135° + isin135°)

OpenStudy (ttop0816):

@hartnn !!

OpenStudy (ttop0816):

is this right?? sqrt2(cos45° + isin45°)

hartnn (hartnn):

nopes

hartnn (hartnn):

sqrt 2 is correct not the angle

hartnn (hartnn):

\(\arctan (\dfrac{1}{-1}) = \arctan(-1) = ... ? \)

OpenStudy (ttop0816):

-pi/4!

hartnn (hartnn):

yes, -45 or -45 + 180

hartnn (hartnn):

as we need a positive angle

OpenStudy (ttop0816):

135! its the second answer (:

hartnn (hartnn):

yep :)

OpenStudy (ttop0816):

this is the last question i promise!! Given the following triangle, if c = sqrt5 and a = 2, find the measure of B to the nearest degree. https://media.glynlyon.com/g_trg_2013/1/rt_triangle_abc.gif B =?? Identify the conic whose equation is given. r=1/3+2sintheta ellipse parabola << right?? hyperbola

hartnn (hartnn):

is that equation correct ? does not seem to be any conic http://www.wolframalpha.com/input/?i=r%3D1%2F3%2B2sintheta

OpenStudy (ttop0816):

r=(1)/(3+2sintheta) (:

hartnn (hartnn):

use cos B = a/c

hartnn (hartnn):

http://www.wolframalpha.com/input/?i=r%3D1%2F%283%2B2sintheta%29 its an ellipse :)

OpenStudy (ttop0816):

oh! thank you((((: then back to the first question! i use cosB= a/c?

hartnn (hartnn):

yes cos angle = adjacent side/ hypotenuse angle here is B

OpenStudy (ttop0816):

cos^-1(2/√5)??

OpenStudy (ttop0816):

cos^1(2/sqrt5)?? then it would be 27 degrees??

hartnn (hartnn):

26.57 yes :)

OpenStudy (ttop0816):

thank you sosososso much for replying back to every SINGLE one of my questions & helping me throughout the whole process! Now im finished with my review & it feels like i've learned waaaay more than school! I think i could do well on my test soon (: thank you again!

hartnn (hartnn):

good luck for the test! :) welcome ^_^